Is my simplifying of this equation correct?
(x+2)^2/x^2+3x+2 = 2/3x

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

Is my simplifying of this equation correct?
(x+2)^2/x^2+3x+2 = 2/3x
-------------------
Sub a value for x
eg, x =5
--------
(5+2)^2/(25 + 15 + 12) =? 2/15
49/52 =? 2/15
Apparently not.

Answer by richwmiller(17219)

(x+2)^2/x^2+3x+2 is not an equation but an expression.
It doesn't simplify to (2/3)*x but to (2+x)/(1+x)